# Multi-Agent Inverted Transformer for Flight Trajectory Prediction

### v1

@@ V Conclusion
Nevertheless, despite its strong performance, several aspects require further exploration to improve its practical applicability in real-world air traffic operations. While MAIFormer provides insights into aircraft interactions through its agent attention, it remains uncertain whether these learned patterns truly reflect a human ATC’s attention. Future research should involve human-in-the-loop simulations or expert assessments to evaluate the interpretability and operational alignment of the attention distributions produced by the proposed method. Furthermore, the model relies solely on flight trajectory data, but in real-world operations, external contextual factors such as meteorological conditions, flight procedures, and airspace constraints play a significant role in ATC decision making and aircraft behavior. Incorporating such conditional information into the model would improve prediction accuracy and bring the model closer to the standard required for real-world deployment in an ATM system.

### v2

@@ V Conclusion
As a potential avenue for practical deployment and future extension, the learned agent-level attention scores can be valuable for a broader range of ATC-related applications. For example, in the development of autonomous ATC systems that aim to imitate historical controller behavior, the aircraft-level attention scores generated by the system can provide insights into how the systems prioritize and monitor specific aircraft in dynamic traffic conditions.
Nevertheless, despite its strong performance, several aspects require further exploration to improve its practical applicability in real-world air traffic operations. MAIFormer provides insights into aircraft interactions through its agent attention, but it remains uncertain whether these learned patterns truly reflect a human ATC’s attention. Future research should involve human-in-the-loop simulations or expert assessments to evaluate the interpretability and operational alignment of the attention distributions produced by the proposed method. Furthermore, the model relies solely on flight trajectory data, but in real-world operations, external contextual factors play a significant role in ATC decision making and aircraft behavior, including meteorological conditions, flight procedures, and airspace constraints. Incorporating such conditional information into the model as exogenous inputs would improve the prediction accuracy and bring the model closer to the standard required for real-world deployment in an ATM system.